# A Computationally Efficient Approach to False Discovery Rate Control and Power Maximisation via Randomisation and Mirror Statistic

### v1

@@ Abstract
Simultaneously performing variable selection and inference in high-dimensional regression models is an open challenge in statistics and machine learning. The increasing availability of vast amounts of variables requires the adoption of specific statistical procedures to accurately select the most important predictors in a high-dimensional space, while controlling the False Discovery Rate (FDR) arising from the underlying multiple hypothesis testing. In this paper we propose the joint adoption of the Mirror Statistic approach to FDR control, coupled with outcome randomisation to maximise the statistical power of the variable selection procedure. Through extensive simulations we show how our proposed strategy allows to combine the benefits of the two techniques. The Mirror Statistic is a flexible method to control FDR, which only requires mild model assumptions, but requires two sets of independent regression coefficient estimates, usually obtained after splitting the original dataset. Outcome randomisation is an alternative to Data Splitting, that allows to generate two independent outcomes, which can then be used to estimate the coefficients that go into the construction of the Mirror Statistic. The combination of these two approaches provides increased testing power in a number of scenarios, such as highly correlated covariates and high percentages of active variables. Moreover, it is scalable to very high-dimensional problems, since the algorithm has a low memory footprint and only requires a single run on the full dataset, as opposed to iterative alternatives such as Multiple Data Splitting.
@@ 1 Introduction
Advances in data collection capabilities have allowed researchers to get access to thousands of features on multiple subjects in relatively short times. Examples are next-generation sequencing technologies that allow fast DNA and RNA sequencing and Nuclear Magnetic Resonance Spectroscopy used to identify metabolites. But also wearable personal devices that can measure several anthropometric variables and clinical outcomes of interests, such as blood pressure and blood glucose levels, continuously over time, thus producing a huge amount of fine-grained measurements. Moreover, it is now common to integrate these multiple input sources into a single study, and repeating these measurements over multiple time points, in a longitudinal fashion, in order to capture potentially relevant trends in the outcomes of interest, given some specific interventions or treatments.
@@
The combination of all these factors generates an incredibly vast and complex set of features that researchers need to analyse with the goal of extracting and estimating the effects of only the most critical drivers related to the biological mechanism under study. This can be a daunting task, for at least two reasons: 1)1)1 ) the available sample size, e.g. the number of patients recruited in a cohort study, is very often much lower than the number of available features (high-dimensional problem). Therefore, standard statistical or machine learning models struggle to extract the underlying fundamental associations, due to model limitations. 2)2)2 ) even when a model can deal with a sample size lower than the number of features, the high probability of selecting false positive effects, caused by the underlying multiple hypothesis testing process, poses a serious threat to the validity of the analysis.
@@
We address the first problem of variable selection by using all available features during the training process, building a statistical model that can automatically select the most relevant variables. This allows to retain the full interpretability about the effects of the covariates on the outcome of interest, in contrast with dimensionality reduction techniques such as Principal Component Analysis [AHT20]. In high-dimensional regression (or classification) problems, popular choices are the LASSO [Tib96], ElasticNet [ZH05], LARS [Efr+04] and SCAD [FL11], which provide efficient algorithms that scale well to a large number of features. Alternatives also exist in the Bayesian framework, where variable selection is performed through an appropriate choice of the prior distributions [OS09]. Popular choices are the two-group Spike and Slab prior distribution, which explicitly provide posterior probabilities of inclusion for each variable; and the one-group prior, such as the Laplace prior distribution, which shrinks coefficients toward zero, acting as a regularisation in a similar fashion to the $L_{1}$ penalty in the LASSO.
@@
In addition to selecting only a subset of variables, in many applications to real data it is also essential to be able to make proper inference on the selected subset of features. This is essential to have a reliable estimate of the regression coefficients confidence intervals and to be able to control some form of error rate, such as the False Discovery Rate (FDR, [BH95]). However, simply proceeding with inference, after a data-dependent variable selection step, does not allow to perform valid post-selection inference, as explained in [Ber+13]. This is because data-driven variable selection procedures, such as the ones mentioned above, generate a model that is not deterministic and the straightforward application of classic approaches to inference, like Ordinary Least Squares, do not account for this additional randomness. Therefore, the estimated coefficients, and the corresponding confidence intervals will be biased, leading to a potential increase in erroneous classifications.
@@
A number of methods exist to control error rates in multiple testing. [MB10] propose stability selection, as a method to control the number of False Discoveries in a LASSO regression setting. [Ber+13] takes a different approach, providing an analytical solution for the linear regression problem when using a LASSO penalty, formally accounting for the variable selection process at the inferential step, thus obtaining correct confidence intervals for the selected coefficients. [RBG22] provide a partial extension to the case of additive and linear mixed models, using Monte Carlo approximations, however the proposed solution is computationally intensive. Overall, this formal treatment is limited to simple cases, such as linear regression, and extensions are difficult.
@@
[BC15], in their pioneering work, propose the knockoff filters procedure as a direct way to control the False Discovery Rate (FDR). The knockoff method works by augmenting the space of covariates, adding a perturbed version of each feature to the design matrix and performing variable selection on the new feature space. The authors provide an upper-bound estimate of the FDR and a new knockoff test statistic through which it is possible to control the FDR at any specific level. This approach, however, has some limitations, i.e. the knowledge of the joint distribution of the covariates is required to construct the knockoff filters and the procedure is limited to the case $n>p$. [Can+18] provide an extension to the high-dimensional scenarios, but still requires knowledge of the joint distribution of the covariates, which is not known in most real data applications.
@@
Despite the limitations of the knockoff, the idea of feature perturbation has been successfully adopted in other works as a way to control FDR. [XZL21] develop the Gaussian Mirrors procedure which allows to control FDR without requiring any distributional assumption on the covariates. However, this approach is inefficient because it only evaluates one variable at a time. Using a similar approach, [Dai+22] develop the Mirror Statistic, by substituting the feature perturbation step with a two-steps procedure based on data splitting (DS). The first step being variable selection on a subset of the data and the second step being statistical inference on a second subset of data, independent of the first. This strategy allows to perform inference simultaneously on all variables and is therefore more efficient. One drawback of this approach is the loss of power caused by the reduced sample size available for the inference step (and similarly for the variable selection step). To mitigate this problem the authors propose a variation called Multiple Data Splitting (MDS), more akin to stability selection, showing increased power in simulations. Nevertheless, MDS is computationally much more costly, since the same procedure has to be repeated multiple times (at least $50$ according to the authors).
@@
In this paper we propose to use the Mirror Statistic, but, instead of creating two independent sub-samples using data splitting, we borrow the idea of outcome randomisation from [TT18], where the authors propose a simple mechanism to create two independent sets of data by adding some random noise to the outcome, splitting more efficiently the original information available into two independent new pseudo-outcomes. The result is an increase in statistical power and a more computationally efficient algorithm. We provide a performance comparison via numerical experiments, replicating the results of [Dai+22] and extending the simulation scenarios to more challenging settings.
@@ 2.1 False Discovery Rate control
The False Discovery Rate has been introduced by [BH95] as a less conservative approach to False Positive error control compared to the Family-Wise Error Rate, which can be too restrictive when testing a large number of hypothesis. FDR is defined as the expectation of the False Discovery Proportion, defined as:
@@
[BH95] provides a correction method for the p-values so that a specific FDR level $\alpha$ can be achieved, under the assumption of independence of the p-values. Although some extensions exist that allow for some form of dependence [BY01], for many of the aforementioned algorithms in Section 1, p-values are not available at all (e.g. LASSO). Hence, the necessity of using methods that can achieve FDR control without explicitly calculating p-values, such as the Mirror Statistic (Eq. 2).
@@ 2.2 Data Splitting and Randomisation
The practice of splitting a given dataset into multiple smaller independent slices is common and the standard in most machine learning applications [Bis06], where, generally, a training, validation and testing set are generated from the original sample. While in machine learning, splitting is done to validate the prediction accuracy of a model, in classic statistical inference the same idea can be used to create valid inferential procedures. [Dai+22] use this strategy in order to control FDR via the test statistic Mirror Statistic, defined for variable $\text{X}_{j}$ as:
@@
where $f(|x|,|y|)$ is a non-negative and monotonically increasing function and $\hat{\beta}^{(1)}_{j}$ and $\hat{\beta}^{(2)}_{j}$ are two distinct estimates of the regression coefficient obtained on two independent subsets of the sample. The logic behind Eq. 2 is that for features that are relevant, the corresponding $M_{j}$ will get a positive relatively large value, because the two independent estimates, $\hat{\beta}^{(1)}_{j}$ and $\hat{\beta}^{(2)}_{j}$, will have concordant signs. Conversely, if the estimated coefficients have discordant signs, $M_{j}$ will always get a negative value and if the estimates are relatively small, meaning that probably the features are not relevant, then $M_{j}$ will be small as well.
@@
#$\#\left\{j\in S_{0}:M_{j}>t\right\}\approx\#\left\{j\in S_{0}:M_{j}<-t\right\}\leq\#\left\{j:M_{j}<-t\right\},\forall t>0$ (3)
which can be directly used to approximate the False Discovery Proportion in Eq. 1.
As we can see from the definition, to use the Mirror Statistic estimator we need two independent sets of observations. Although data splitting is universally valid and straightforward to use, it comes at the cost of a much reduced sample size, which can have detrimental effects on the power of the statistical test and on the stability of the variable selection. To counteract this downside [Dai+22] propose MDS as a way to increase the power of the test statistic. MDS amounts to repeating multiple times the whole procedure of variable selection with simple DS and then aggregating the results. In simulations MDS seems to provide higher power; however, this improvement comes at the cost of a much higher computational burden and an additional uncertainty due to the choice of the aggregation strategy.
This is where Randomisation [TT18] offers an alternative way of distributing the available sample information and helps avoiding the randomness of the simple data splitting, creating two pseudo-independent sets, by perturbing the outcome with some random noise $W$. The general idea is to use a randomisation scheme through which we only allow ourselves to observe the outcome through $U\equiv u(Y,W)$ at the variable selection step, while at the inference step we only observe $V\equiv v(Y,W)$, with $V$ constructed to be independent of $U$.
@@
Given $\sigma^{2}$, or an estimate $\hat{\sigma}^{2}$, we can generate an $n$-dimensional vector of random Normal noise as $\boldsymbol{w}\sim\text{N}_{n}(\boldsymbol{0},\sigma^{2}\gamma\text{I}_{n})$, where the scalar $\gamma>0$ allows to distribute information for variable selection and inference ($\gamma=1$ corresponds to splitting data in half, while $\gamma<1$ places more information on $U$). Then, building $U=Y+W$, we have that $\boldsymbol{u}\sim\text{N}_{n}(\boldsymbol{\mu},\sigma^{2}(1+\gamma)\text{I}_{n})$ and $V=Y-\gamma^{-1}W$, with $\boldsymbol{v}\sim\text{N}_{n}(\boldsymbol{\mu},\sigma^{2}(1+\gamma^{-1})\text{I}_{n})$.
Randomisation can be interpreted as averaging information over all possible data splits of the same size. [TT18] show that for a Normally distributed outcome $\boldsymbol{y}$, randomisation guarantees a power that is always at least as high as data-splitting.
@@
$\gamma$, $q$ (FDR target level), $\hat{\sigma}^{2}$
$\boldsymbol{w}\overset{IID}{\sim}\text{N}_{n}(\boldsymbol{0},\sigma^{2}\gamma\text{I}_{n})$
$\boldsymbol{u}=\boldsymbol{y}+\boldsymbol{w}$
$\boldsymbol{v}=\boldsymbol{y}-\gamma^{-1}\boldsymbol{w}$
Estimate regression coefficients $\boldsymbol{\hat{\beta}^{(u)}}$ and $\boldsymbol{\hat{\beta}^{(v)}}$ from $\boldsymbol{u}$ and $\boldsymbol{v}$ respectively.
Calculate MS $M_{j}$, $\forall j$, using Eq. 2.
Select the covariates $\text{X}_{j}$ for which ${M_{j}>\tau_{q}}$, where the threshold $\tau_{q}$ is

### v2

@@ Abstract
Simultaneously performing variable selection and inference in high-dimensional regression models is an open challenge in statistics and machine learning. The increasing availability of vast amounts of variables requires the adoption of specific statistical procedures to accurately select the most important predictors in a high-dimensional space, while controlling the False Discovery Rate (FDR) associated with the variable selection procedure. In this paper we propose the joint adoption of the Mirror Statistic approach to FDR control, coupled with outcome randomisation to maximise the statistical power of the variable selection procedure, measured through the True Positive Rate (TPR). Through extensive simulations we show how our proposed strategy allows to combine the benefits of the two techniques. The Mirror Statistic is a flexible method to control FDR, which only requires mild model assumptions, but requires two sets of independent regression coefficient estimates, usually obtained after splitting the original dataset. Outcome randomisation is an alternative to Data Splitting, that allows to generate two independent outcomes, which can then be used to estimate the coefficients that go into the construction of the Mirror Statistic. The combination of these two approaches provides increased testing power in a number of scenarios, such as highly correlated covariates and high percentages of active variables. Moreover, it is scalable to very high-dimensional problems, since the algorithm has a low memory footprint and only requires a single run on the full dataset, as opposed to iterative alternatives such as Multiple Data Splitting.
keywords:
false discovery rate, gene expression, high-dimensional regression, post-selection inference, variable selection
@@ 1 Introduction
Advances in data collection capabilities have allowed researchers to get access to thousands of features on multiple subjects in relatively short times. Examples are next-generation sequencing technologies that allow fast DNA and RNA sequencing and Nuclear Magnetic Resonance Spectroscopy used to identify metabolites. 1 But also wearable personal devices that can measure several anthropometric variables and clinical outcomes of interests, such as blood pressure and blood glucose levels, continuously over time, thus producing a vast collection of measurements. 2 Moreover, it is now common to integrate these multiple input sources into a single study, and repeating these measurements over multiple time points, in a longitudinal fashion, in order to capture potentially relevant trends in the outcomes of interest, given some specific interventions or treatments.
@@
The combination of all these factors generates an incredibly vast and complex set of features that need to be jointly analyzed. 3; 4 In many such studies, feature selection becomes an important task, a typical example being biomarker discovery. This can be a daunting task, for at least two reasons: 1)1)1 ) the available sample size, e.g. the number of patients recruited in a cohort study, is very often much lower than the number of available features (high-dimensional problem). Therefore, standard statistical or machine learning models struggle to extract the underlying fundamental associations, due to model limitations. 2)2)2 ) even when a model can deal with a sample size lower than the number of features, the high probability of selecting false positive effects, poses a serious threat to the validity of the analysis.
@@
We address the first problem of variable selection by using all available features during the training process, building a statistical model that can automatically select the most relevant variables. This allows to retain the full interpretability about the effects of the covariates on the outcome of interest, in contrast with dimensionality reduction techniques such as Principal Component Analysis.5 In high-dimensional regression (or classification) problems, popular choices are the LASSO, ElasticNet, LARS and SCAD, which provide efficient algorithms that scale well to a large number of features.6; 7; 8; 9 Alternatives also exist in the Bayesian framework, where variable selection is performed through an appropriate choice of the prior distributions.10 Popular choices are the two-group Spike and Slab prior distribution, which explicitly provide posterior probabilities of inclusion for each variable; and the one-group prior, such as the Laplace prior distribution, which shrinks coefficients toward zero, acting as a regularisation in a similar fashion to the $L_{1}$ penalty in the LASSO.
@@
In addition to selecting only a subset of variables, in many applications to real data it is also essential to be able to make proper inference on the selected subset of features. This is essential to have a reliable estimate of the regression coefficients confidence intervals and to be able to control some form of error rate, such as the False Discovery Rate (FDR).11 However, simply proceeding with inference, after a data-dependent variable selection step, does not allow to perform valid post-selection inference, as explained in Berk et al.12 This is because data-driven variable selection procedures, such as the ones mentioned above, generate a model that is not deterministic and the straightforward application of classic approaches to inference, like Ordinary Least Squares, do not account for this additional randomness. Therefore, the estimated coefficients, and the corresponding confidence intervals will be biased, leading to a potential increase in erroneous classifications.
@@
A number of methods exist to control error rates in multiple testing. Meinshausen and Buhlmann13 propose stability selection, as a method to control the number of False Discoveries in a LASSO regression setting. Lee et al.14 takes a different approach, providing an analytical solution for the linear regression problem when using a LASSO penalty, formally accounting for the variable selection process at the inferential step, thus obtaining correct confidence intervals for the selected coefficients. Rugamer, Baumann and Greven15 provide a partial extension to the case of additive and linear mixed models, using Monte Carlo approximations, however the proposed solution is computationally intensive. Overall, this formal treatment is limited to simple cases, such as linear regression, and extensions are difficult.
@@
Barber and Candes,16 in their pioneering work, propose the knockoff filters procedure as a direct way to control the False Discovery Rate (FDR). The knockoff method works by augmenting the space of covariates, adding a perturbed version of each feature to the design matrix and performing variable selection on the new feature space. The authors provide an upper-bound estimate of the FDR and a new knockoff test statistic through which it is possible to control the FDR at any specific level. This approach, however, has some limitations, i.e. the knowledge of the joint distribution of the covariates is required to construct the knockoff filters and the procedure is limited to the case $n>p$. Candes et al.17 provide an extension to the high-dimensional scenarios, but still requires knowledge of the joint distribution of the covariates, which is not known in most real data applications.
@@
Despite the limitations of the knockoff, the idea of feature perturbation has been successfully adopted in other works as a way to control FDR. Xing, Zhao and Liu18 develop the Gaussian Mirrors procedure which allows to control FDR without requiring any distributional assumption on the covariates. However, this approach is inefficient because it only evaluates one variable at a time. Using a similar approach, Dai et al.19 develop the Mirror Statistic, by substituting the feature perturbation step with a two-steps procedure based on data splitting (DS). The first step being variable selection on a subset of the data and the second step being statistical inference on a second subset of data, independent of the first. Since the two sets are independent by construction, the conditional distribution of the inference set given the output of the selection step is the same as the unconditional one, so classical procedures can be used to provide valid inference for the selected parameters.20 One drawback of this approach is the loss of power caused by the reduced sample size available for the inference step (and similarly for the variable selection step). To mitigate this problem the authors propose a variation called Multiple Data Splitting (MDS), more akin to stability selection, showing increased power in simulations.19 The authors show that this approach outperforms the knockoff filter in many situations. Nevertheless, MDS is computationally much more costly than DS, since the same procedure has to be repeated multiple times (at least $50$ according to the authors).
@@
In this paper we propose to use the Mirror Statistic, but, instead of creating two independent sub-samples using data splitting, we borrow the idea of outcome randomisation from Rasines and Young20, where the authors propose a simple mechanism to create two independent sets of data by adding some random noise to the outcome, splitting more efficiently the original information available into two independent new pseudo-outcomes. The result is an increase in statistical power and a more computationally efficient algorithm. We provide a performance comparison via numerical experiments, replicating the results of Dai et al.19 and extending the simulation scenarios to more challenging settings.
@@ 2.1 False Discovery Rate control
The False Discovery Rate has been introduced as a less conservative approach to False Positive error control compared to the Family-Wise Error Rate, which can be too restrictive when testing a large number of hypothesis.11 FDR is defined as the expectation of the False Discovery Proportion:
@@
Benjamini and Hochberg11 provide a correction method for the p-values so that a specific FDR level $\alpha$ can be achieved, under the assumption of independence of the p-values. Although some extensions exist that allow for some form of dependence,21 for many of the aforementioned algorithms in Section 1, p-values are not available at all (e.g. LASSO). Hence, the necessity of using methods that can achieve FDR control without explicitly calculating p-values, such as the Mirror Statistic (Equation 2).
For completeness we also specify the formula for the True Positive Rate (TPR), defined as
$\text{TPR}=\dfrac{\#\left\{j:j\in S_{1},j\in\hat{S}\right\}}{\#\left\{j:j\in S_{1}\right\}}$
The TPR is used to measure the power of the method, i.e. the ability to select the true active variables.
@@ 2.2 Data Splitting and Randomisation
The practice of splitting a given dataset into multiple smaller independent slices is common and the standard in most machine learning applications, 22 where, generally, a training, validation and testing set are generated from the original sample. While in machine learning, splitting is done to validate the prediction accuracy of a model, in classic statistical inference the same idea can be used to create valid inferential procedures. Dai et al.19 use this strategy in order to control FDR via the test statistic Mirror Statistic, defined for variable $\text{X}_{j}$ as:
@@
where $f(|x|,|y|)$ is a non-negative, exchangeable and monotonically increasing function and $\hat{\beta}^{(1)}_{j}$ and $\hat{\beta}^{(2)}_{j}$ are two distinct estimates of the regression coefficient obtained on two independent subsets of the sample. The logic behind Equation 2 is that for features that are relevant, the corresponding $M_{j}$ will get a positive relatively large value, because the two independent estimates, $\hat{\beta}^{(1)}_{j}$ and $\hat{\beta}^{(2)}_{j}$, will have concordant signs. Conversely, if the estimated coefficients have discordant signs, $M_{j}$ will always get a negative value and if the estimates are relatively small, meaning that probably the feature is not relevant, then $M_{j}$ will be small as well.
@@
#$\begin{split}&\#\left\{j\in S_{0}:M_{j}>t\right\}\\ \approx\ &\#\left\{j\in S_{0}:M_{j}<-t\right\}\\ \leq\ &\#\left\{j:M_{j}<-t\right\},\forall t>0\end{split}$ (3)
which can be directly used to approximate the False Discovery Proportion in Equation 1.
As we can see from the definition, to use the Mirror Statistic estimator we need two independent sets of observations. Although data splitting is universally valid and straightforward to use, it comes at the cost of a much reduced sample size, which can have detrimental effects on the power of the statistical test and on the stability of the variable selection. To counteract this downside Dai et al.19 propose MDS as a way to increase the power of the test statistic. MDS amounts to repeating multiple times the whole procedure of variable selection with simple DS and then aggregating the results. In simulations MDS seems to provide higher power; however, this improvement comes at the cost of a much higher computational burden and an additional uncertainty due to the choice of the aggregation strategy.
This is where Randomisation offers an alternative way of distributing the available sample information and helps avoiding the randomness of the simple data splitting, creating two pseudo-independent sets, by perturbing the outcome with some random noise $W$20 The general idea is to use a randomisation scheme through which we only allow ourselves to observe the outcome through $U\equiv u(Y,W)$ at the variable selection step, while at the inference step we only observe $V\equiv v(Y,W)$, with $V$ constructed to be independent of $U$.
@@
Given $\sigma^{2}$, or an estimate $\hat{\sigma}^{2}$, we can generate an $n$-dimensional vector of random Normal noise as $\boldsymbol{w}\sim\text{N}_{n}(\boldsymbol{0},\sigma^{2}\gamma\text{I}_{n})$, where the scalar $\gamma>0$ allows to distribute information for variable selection and inference, respectively. Rasines and Young20 show how using $\gamma=1$ is equivalent to splitting the data in two halves of equal size, which is the most natural choice in absence of any additional information. Any other choice would place more information on either the LASSO or the OLS.
Then, building $U=Y+W$, we have that $\boldsymbol{u}\sim\text{N}_{n}(\boldsymbol{\mu},\sigma^{2}(1+\gamma)\text{I}_{n})$ and $V=Y-\gamma^{-1}W$, with $\boldsymbol{v}\sim\text{N}_{n}(\boldsymbol{\mu},\sigma^{2}(1+\gamma^{-1})\text{I}_{n})$.
Randomisation can be interpreted as averaging information over all possible data splits of the same size. Rasines and Young20 show that for a Normally distributed outcome $\boldsymbol{y}$, randomisation guarantees a power that is always at least as high as data-splitting.
@@
1:$\gamma$, $q$ (FDR target level), ${\sigma}^{2}$
2:$\boldsymbol{w}\overset{IID}{\sim}\text{N}_{n}(\boldsymbol{0},\sigma^{2}\gamma\text{I}_{n})$
3:$\boldsymbol{u}=\boldsymbol{y}+\boldsymbol{w}$
4:$\boldsymbol{v}=\boldsymbol{y}-\gamma^{-1}\boldsymbol{w}$
5:Estimate regression coefficients $\boldsymbol{\hat{\beta}^{(u)}}$ and $\boldsymbol{\hat{\beta}^{(v)}}$ from $\boldsymbol{u}$ and $\boldsymbol{v}$ respectively.
6:Calculate MS $M_{j}$, $\forall j$, using Equation 2.
7:Select the covariates $\text{X}_{j}$ for which ${M_{j}>\tau_{q}}$, where the threshold $\tau_{q}$ is
@@
If the residual variance ${\sigma}^{2}$ is unknown, it can be estimated using the residual sum of squares obtained from the LASSO model fitted with penalisation parameter tuned by $k$-fold cross-validation,20, that is:
$\hat{\sigma}^{2}=\sum_{i=1}^{\tilde{n}}\dfrac{\left(y_{i}-\hat{y}_{i}\right)^{2}}{\tilde{n}-df-1}$
where $\tilde{n}$ is the sample size, $df$ is the number of non-zero coefficients and $\hat{y}$ is the predicted outcome.
This is the strategy that we adopt in the numerical experiments in Section 3, as well as in the real-world data application in Section 4, where in both cases we use $k=10$ folds.
We further explore the implications of an incorrectly estimated variance ${\sigma}^{2}$ through additional simulations where we control the value of ${\sigma}^{2}$ used in the model. In the Appendix, we report the results of this experiment. When the variance is underestimated the FDR is not properly controlled, while when the variance is overestimated the algorithm becomes more conservative in terms of FDR control.
2.3 Assumptions
The theory underlying the development of the Mirror Statistic is based on the independence of the two sets of coefficient estimates and the following symmetry and weak dependence assumptions:19
  1. 1.
    Symmetry: For each feature $j\in S_{0}$, the sampling distribution of at least one of the two coefficients, $\hat{\beta}^{(1)}_{j}$ and $\hat{\beta}^{(2)}_{j}$, is symmetric around zero
  2. 2.
    Weak dependence: The mirror statistics $M_{j}$ are continuous random variables, and there exist a constant $c>0$ and $\alpha\in(0,2)$ such that
    $\text{Var}\left(\#\left\{M_{j}>t\mid j\in S_{0}\right\}\right)\leq cp_{0}^{\alpha},\forall t>0$
    where $p_{0}$ is the number of null features. This assumption translates into a restriction on the correlation among the null features
A requirement for the symmetry condition is that all variables with a non-null coefficient are selected in the variable selection step. This is generally referred to as the sure screening property. In our simulations and real-world application we use the combination of LASSO and OLS, respectively to perform variable selection and inference, in linear models. For the LASSO, the sure screening property depends on the Signal strength condition, defined as:
Signal strength $\textit{Signal strength condition}:\underset{j\in S_{1}}{\text{min}}|\beta_{j}|\gg\sqrt{p_{1}\log p/n}$
In order to better understand the impact of violating the assumption of symmetry we run a controlled simulation where we monitor the individual steps of the randomization plus Mirror Statistic algorithm. Results are detailed in the Appendix.